Project

General

Profile

Run Plan » History » Version 268

« Previous - Version 268/754 (diff) - Next » - Current version
Andrew Olivier, 06/22/2015 11:32 AM


Run Plan

Please help keep this list complete and up to date.

Known alarms

It is ok to ignore the following alarms in Synoptic:

  • MuRS 12 Voltage: This PMT is intentionally turned off off until the paddles are replaced.

Upcoming beam-off periods

- Next LAr delivery : Wednesday 2015.06.24 7:30 am. Open access and gate re-closure will be arranged.
- Measure muon range stack thicknesses and confirm placement of small versus large cosmic paddles
- 2015.06.17 6am Roughly one day -- Check with MCR. Down for Booster RF Cavity replacement:
- 2015.06.18 7am - 9am and briefly at 11am -- Preparation in Switchyard for Main Injector Beam-On study

Beamy Priorities

Current Runs

Beam Request Tertiary Beam Magnet Polarity/Current Trigger Configuration Duration Notes
+32 GeV, 100,000 @ SC1 +100 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4-MURS-PUNCH
Beginning 2015.06.20 Adjust intensity at SC1 to get as many few-track events as possible, in the TPC.
-32 GeV, 100,000 @ SC1 -100 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4-MURS-PUNCH
Beginning 2015.06.16 Adjust intensity at SC1 to get as many few-track events as possible, in the TPC.
-16 GeV, 18,000 @ SC1 -100 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4-MURS-PUNCH
Days (Beginning 2015.05.30) Done for now. Adjust intensity at SC1 to get as many few-track events as possible, in the TPC.
-16 GeV, 55,000 @ SC1 -60 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-PUNCH-MURS
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-MURS
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-PUNCH
BEAMON+USTOF+DSTOF+WCCOINC3OF4-MURS-PUNCH
Days (Beginning 2015.06.09) Done for now Adjust intensity at SC1 to get as many few-track events as possible, in the TPC.

Run table

Beam Request Tertiary Beam Magnet Polarity/Current Trigger Configuration Duration Notes
60 GeV, 30,000 @ SC1 100 A Trigger0: on:(BEAMON USTOF DSTOF WCCOINC3OF4) off:(HALO),
Trigger1: on: (COSMICON COSMIC)
Days Physics data Complete

Beamless Priorities

Tasks

  • Tangle-by-the-door:
    • Return cables to FTBF.
    • Render unto PREP that which is PREP's.
  • Connect COW2 to ACNET. Same correction scheme needed for reported values?

Empty-event runs

to measure the RMS of the noise at each Gain & Filter setting.

Only one trigger pattern: on: PULSER. Take a 3-spill run at each setting combination and then ensure that the first spill appears shows Data Blocks from the V1740.

Record the first spill and event with a confirmed-empty event display:
Gain 3, Filter 0
r5636 sp1 ev0
Gain 3, Filter 1
r5635 sp1 ev0
Gain 3, Filter 2
r5528 sp1 ev0
Gain 3, Filter 3
r5526 sp1 ev5
Gain 2, Filter 0
r5637 sp1 ev3
Gain 2, Filter 1
r5638 sp1 ev0
Gain 2, Filter 2
r5529 sp1 ev0
Gain 2, Filter 3
r5518 sp1 ev119
Gain 1, Filter 0
r5640 sp1 ev0
Gain 1, Filter 1
r5639 sp1 ev0
Gain 1, Filter 2
r5532 sp4 ev9
Gain1, Filter 3
r5520 sp1 ev0
Gain 0, Filter 0
r5542 sp1 ev1
Gain 0, Filter 1
r5538 sp1 ev1
Gain 0, Filter 2
r5641 sp2 ev0
Gain0, Filter 3
r5527 sp1 ev0

Purity measurement

Needs some really good tracks, hopefully at least one good, straight Cathode-to-Anode track. Take them during the COSMICON window with these trigger patterns:

Beam:
  • 60 GeV on target, 30,000 counts at SC1
Trigger Pattern:
  • Trigger0: on:(BEAMON USTOF DSTOF WCCOINC3OF4) off:(HALO)
  • Trigger1: on: (COSMICON COSMIC)
Record any nice cosmic events (straight cathode-to-anode tracks) here: Cathode is at ~2800 ticks
Gain 3, Filter 0
r sp ev
Gain 3, Filter 1
r sp ev
Gain 3, Filter 2
r sp ev
Gain3, Filter 3
r5435 sp3 ev81
r5605 sp38 ev6
r5605 sp250 ev2
r5605 sp341 ev4
r5605 sp369 ev1
r5605 sp374 ev2
r5834 sp11 ev2
r5854 sp283 ev0
Gain 2, Filter 0
r sp ev
Gain 2, Filter 1
r sp ev
Gain 2, Filter 2
r sp ev
Gain 2, Filter 3
r5441 sp3 ev3
Gain 1, Filter 0
r sp ev
Gain 1, Filter 1
r sp ev
Gain 1, Filter 2
r sp ev
Gain1, Filter 3
r5479 sp47 ev3
Gain 0, Filter 0
r sp ev
Gain 0, Filter 1
r sp ev
Gain 0, Filter 2
r sp ev
Gain0, Filter 3
r5451 sp8 ev3
r5451 sp17 ev2

Completed Runs Below

Beam Request Tertiary Beam Magnet Polarity/Current Trigger Configuration Duration Notes
+16 GeV, 18,000 @ SC1 -60 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-HALO
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-HALO
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-HALO
Days Starting 2015.05.30
+16 GeV, 18,000 @ SC1 100 A BEAMON+USTOF+DSTOF+WCCOINC3OF4-HALO
COSMICON+COSMIC
BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH-HALO
BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS-HALO
Days Done 2015.05.30
-16 GeV, 15,000 @ SC1 -100 A BEAMON+USTOF+DSTOF+WCCOINC3OF4 COSMICON+COSMIC BEAMON+USTOF+DSTOF+WCCOINC3OF4+PUNCH BEAMON+USTOF+DSTOF+WCCOINC3OF4+MURS COSMICON+MICHEL MICHEL-BEAMON Veto: Days (Beginning 2015.06.05 from run 6042) 30-50 spills each. ETL PMT voltage changing from -1257 going down (up in abs value, to -1420V). v1751 readout time extended from ~14us to ~28 us
16 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON USTOF DSTOF WCCOINC3OF4)
Trigger1: on: (COSMICON COSMIC)
Establish trigger rate per SC5 counts. ~Few spills. Measure rates wrt SC5 Done
16 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON USTOF DSTOF WCCOINC3OF4) off: (HALO)
Trigger1: on: (COSMICON COSMIC)
enough MWPC triggers to establish a baseline (~100k) Measure rates wrt SC5 Confirm effect of ~HALO. Done
16 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON COSMIC)
Trigger1: on: (BEAMON HALO,
Trigger2: on: (BEAMON USTOF DSTOF WCCOINC3OF4 HALO)
Trigger3: on: (BEAMON USTOF DSTOF WCCOINC3OF4 COSMIC)
Trigger4: on: (BEAMON COSMIC) off: (HALO)
Trigger5: on; (BEAMON COSMIC HALO DSTOF)
Enough to be sure of rate. Measure trigger rates wrt SC5 Done
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on:(BEAMON USTOF DSTOF WCCOINC3OF4)
Trigger1: on: (COSMICON COSMIC)
Establish trigger rate per SC5 counts. ~Few spills. Measure rates wrt SC5 Done
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON USTOF DSTOF WCCOINC3OF4) off: (HALO)
Trigger1: on: (COSMICON COSMIC)
enough MWPC triggers to establish a baseline (~100k) Measure rates wrt SC5 Confirm effect of ~HALO. Done
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON COSMIC)
Trigger1: on: (BEAMON HALO)
Trigger2: on: (BEAMON USTOF DSTOF WCCOINC3OF4 HALO)
Trigger3: on: (BEAMON USTOF DSTOF WCCOINC3OF4 COSMIC)
Trigger4: on: (BEAMON COSMIC) off: (HALO)
Trigger5: on: (BEAMON COSMIC HALO DSTOF)
Enough to be sure. Measure trigger rates wrt SC5 Done
Beam Request Tertiary Beam Magnet Polarity/Current Trigger Configuration Duration Notes
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 0 in the TDC.config. Measure S/N for the WCs. Done
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 511 in the TDC.config. Measure S/N for the WCs. Done r5090
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 1023 in the TDC.config. Measure S/N for the WCs. Done r5088
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 1535 in the TDC.config. Measure S/N for the WCs. Done r5087
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 2047 in the TDC.config. Measure S/N for the WCs. Done r5086
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 2559 in the TDC.config. Measure S/N for the WCs. Done r5085
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 3071 in the TDC.config. Measure S/N for the WCs. Done r5084
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 3583 in the TDC.config. Measure S/N for the WCs. Done r5091
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON SC1 SC2 SC3 USTOF DSTOF)
Trigger1: on: (COSMICON COSMIC)
3k MWPC triggers Changing IBLR to 4095 in the TDC.config. Measure S/N for the WCs. Done r5103
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON COSMIC)
Trigger1: on: (BEAMON HALO)
Trigger2: on: (BEAMON COSMIC) off: (HALO)
Trigger3: on: (BEAMON COSMIC HALO)
Trigger4: on:(BEAMON) off: (COSMIC)
Trigger5: on: (BEAMON) off:(HALO)
Trigger6: on: (BEAMON HALO) off:(COSMIC)
Enough to be sure. (~30 Spills) Measure trigger rates wrt SC5. Does !HALO really raise the rate? Done run 5104
32, 16, 8 GeV pi+, maximizing MC7SC5 +100 A Trigger0: on: (BEAMON USTOF DSTOF WCCOINC3OF4)
Trigger1: on:(BEAMON USTOF DSTOF WCCOINC3OF4) off:(HALO)
Trigger2: on:(BEAMON USTOF DSTOF WCCOINC3OF4) off: (HALO COSMIC)
Trigger3: on: (BEAMON USTOF DSTOF WCCOINC3OF4) off: (COSMIC)
Trigger4: on: (BEAMON HALO)
Trigger5: on: (BEAMON COSMIC)
30 spills at 32 GeV, 60 spills at 16 GeV, enough stats (few days?) @ 8 GeV Determine possible sweet spot/ trend wrt energy Started--- (Done Run 5105 at 32GeV), (Done Run 5106 no data 5129 @16GeV)
Future ideas:
Beam Request Tertiary Beam Magnet Polarity/Current Trigger Configuration Duration Notes
32 GeV pi+, maximizing MC7SC5 +100 A Trigger0: BEAMON USTOF DSTOF WCCOINC3OF4, Trigger1: COSMICON COSMIC enough MWPC triggers to establish a baseline (~100k) Analyze for coincidence with HALO and COSMIC. Occupancy much higher than 8, 16 GeV?
8 GeV pi+, maximizing MC7SC5 +100 A SC1 SC2 SC3 (all Thresh. Discr.) enough MWPC triggers to establish a baseline Starting 2015/04/02 17h, but put on hold.
8 GeV pi+, maximizing MC7SC5 +100 A SC1 SC2 SC3 (all Const. Frac. Discr.) same MWPC trigger count as above
8 GeV pi+, maximizing MC7SC5 -100A SC1 SC2 SC3 WC4 ~10k triggers Occupancy study for negatives
Studies and Tasks
  1. PT paddle 3: Repair Done 2015.04.22 and re-install. (Planned 2015.04.23)
  2. Connect all 16 trigger inputs to the remaining channels of the V1740 bridge.
  3. (ongoing) Aerogel characterization & studies - Requires tracking be added to analysis/DQM, for position dependence.
  4. Try Halo veto in triggers -- first, estimate rate in data, if possible
  5. Using punchthrough & Halo counters, study halo activity in TPC area. Also try the cosmic finders. What do we expect, out of time with the trigger? Measure beam halo vs 2ndary beam parameters (energy, focus, MTest activity…)
  6. Plateau µRS vs HV
  7. Measure µRS penetration histogram

TPC Voltages

Based on e-log entry #2544:
Plane Setting Readback
Shield -298V -298.75V
Induct -27V -18.5V
Collect +340V +338.4V

HV installation

  1. Done 2015.04.13 Run cables from Bias filter box to PP. Bundle is already labeled and rests on the cryostat table. Check for TPC noise changes before/after connection. 04/10/2015
  2. Confirm connectivity, resistivity of SHV system.
  3. Confirm proportional voltages on TPC for drift, grid plane, induction plane, collection plane
  4. ORC 2015.04.14 16h00

Cryogenic Installation

NB All dates assume that no major problems are encountered.

Done done
  1. Foam insulation of the cryogenic lines inside MC7, happening by Rob week of 3/16/15. Done 2015.03.20
  2. The lines need to be connected to the LAr dewar, hopefully done this week. Tim Griffin and Tommy O. are actively working on it. Done 2015.03.20
  3. Labeling/tagging of dewar Done 2015.03.26
  4. ODH system test and connection to FIRUS. Connected, but system test failed (as of 3/31/15). FIRUS programming error fixed, and both systems are live. Done 04/01/2015
  5. Operate and check calibration of both Badger control valves, which includes the Badger at the ceiling. Use the bottled gas cylinders in the enclosure to run valves. Done 04/01/2015
  6. Fix wired control valve missing from the graphics found 04/01/2015. Done 04/07/2015
  7. The leaking heater section needs to be replaced. Skyler is working on producing the new section starting 3/18/15. Awaiting the high-heat epoxy 3/27/15. Epoxy arrived and being installed as of 04/07/15. Estimated installation 04/08/15. Done 04/09
  8. Tube trailer moved Done 3/30 and connected Done 4/9/2015.
  9. Fill LAr dewar, (Thursday 3/26) if okay by Safety Committee. Done 2015.03.26 Installation of pressure and level telemetry equipment by TWC. Connection to the control system. Done before 2014.04.07
  10. The pressure test of the heater section and lines to the LAr dewar with Bob Barger after the heater section is done and Bob is available. Done 4/10/2015
  11. Installation of PT-403-H for monitoring flow out of tube trailer, connection to control system. Done as of 4/10/15
  12. Foam Insulation of new section of fill line between dewar and building, after pressure test, by Rob. Done 4/10/15
  13. Pyrogel insulation of the heater section after the pressure test. Done 4/13/15
  14. Wiring and checking for remote control for one vacuum pump. Done 04/13/2015
  15. Controls check out 4/6-4/8??. Address logic errors found 04/01/2015. Done as of 4/13/15
Tube Trailer
  1. Regeneration of filter. Done
Controls
  1. Calibrate the analyzers once the gas from the dewar is hooked up. Done
System wrap-up and start-up
  1. Cryogenic Safety walk through, Tuesday 4/14/15 @ 8:30 am. Done, Cryo Safety recommendation for running approved 4/14/15.
  2. Resolution of walk thru findings. ORC is granted, 4/14 or 4/15. Done
  3. Start up of gas analyzers, 4/15. Done
  4. Regeneration of Filters, 4/15-4/16. Done
  5. Evacuate and backfill several times 4/17. Done
  6. Cooldown Filter ~ 1 day Done
  7. Cooldown and fill of LArIAT cryostat. Done